If a person's eye is x meters above the sea level, and she can see y meter in the horizon, then y=square root 3.75 square root x. What if the eye level is 5.29 meters above sea level, how far can she see in the horizon?


1
Expert's answer
2022-01-17T16:19:23-0500
"y=\sqrt{3.75}\cdot\sqrt{5.29}=1.15\sqrt{15} (m)"

She can see "1.15\sqrt{15}" meter in the horizon.